OpenStudy (anonymous):

Solve and check: 2/3 + 5/y-4 = y+6/3y-12

OpenStudy (anonymous):

Just wondering, did you mean \[2/3+5/(y-4)=(y+6)/(3y-12)?\]

OpenStudy (anonymous):

2/3+5/y-4=y+6/3y-12 5/(y-4)-[y+6/3y-12]=-2/3 (15-3y-18)/3y-12=-2/3 -3y-3/3y-12=-2/3 -3/3[y+1/y-4]=-2/3 y+1/y-4=2/3 3(y+1)=2(y-4) 3y+3=2y-8 3y-2y=-8-3 y=-11

OpenStudy (anonymous):

yes, and thank u so much